# Second Order Properties of Thinned Counts in Finite Birth–Death Processes

### v1

@@ Abstract
The paper studies the counting process arising as a subset of births and deaths in a birth–death process on a finite state space. Whenever a birth or death occurs, the process is incremented or not depending on the outcome of an independent Bernoulli experiment whose probability is a state-dependent function of the birth and death and also depends on whether it is a birth or death that has occurred. We establish a formula for the asymptotic variance rate of this process, also presented as the ratio of the asymptotic variance and the asymptotic mean. Several examples including queueing models illustrate the scope of applicability of the results. An analogous formula for the countably infinite state space is conjectured and tested.
@@ 1 Introduction
We first became aware of the second-order properties discussed in this note in various queueing models in which the simpler cases can be cast purely in a birth–death process setting where computation is more readily effected. In such cases there is a phenomenon, called BRAVO below, that appears in the limit as the state space grows. In this paper we exhibit more general processes $N_{q}$ defined by counting the births and deaths that remain after independent thinnings of these two types of events in a birth–death process and for which the BRAVO characteristics may or may not survive the thinning procedure.
@@
$\mathcal{D}:=\lim_{t\to\infty}\frac{\mathop{\rm var}\nolimits N_{q}(0,t]}{{\mathbb{E}}\big{[}N_{q}(0,t]\big{]}}.$ (1.3)
In mathematical biology, for a general stationary counting process $N_{q}$, $\mathcal{D}$ is known as the Fano factor, as for example in Eden and Kramer [6]. The denominator and numerator here are asymptotically linear in $t\geq 0$, and the limit in (1.3) is always finite and positive. Our main result, Theorem 1, presents a formula for $\mathcal{D}$ in terms of $\{q^{\pm}_{j}:j\in\mathbb{J}\}$ and the birth- and death-rates of $Q$. To our knowledge this condensed explicit formula is new. Yet we mention that since the state space is finite, computable matrix based expressions for $\mathcal{D}$ can be obtained by considering the process as a MAP (Markovian Arrival Process); see for example [9].
In the special case of a pure death-counting process (all $q_{i}^{+}=0$), no thinning ($q^{-}_{i}\equiv 1$), a similar formula was derived in [14] where the BRAVO effect was first noted. This phenomenon occurs when the birth–death process $Q(t)$ models a queueing system and $N_{-}$ is the output process of served customers. In a sequence of papers associated with BRAVO [4, 5, 7, 8, 11, 12, 13, 14], it was shown that for non-small $J$, it holds that $\mathcal{D}\approx 1$, except for ‘balanced‘ singular cases in which the arrival rate equals the service capacity. In those cases $\mathcal{D}\approx{\cal B}$ where for many standard queueing models, the BRAVO constant ${\cal B}<1$. This constant equals $2/3$ for simple single-server Markovian queues with a finite buffer, [8, 14]; equals $2(1-2/\pi)\approx 0.7268$ for infinite buffer cases with infinite buffers, [1, 13]; and equals $\approx 0.6093$ in many-server queues under the so-called Halfin–Whitt scaling regime, [5]. Note also a correction paper for some of the computations associated with that result, [11].
Further, for queues with renewal input (i.e. GI//$\cdot$/$\cdot$)⋅ / ⋅ ) systems), the first multiple, $2$, in $2(1-2/\pi)$ is replaced by the sum of the squared coefficients of variation of the inter-arrival and service times driving a GI/G/$s$ queue, [1]; similarly in the finite case, as conjectured and numerically tested, [12], ${\cal B}$ equals the sum of the squared coefficients of variations divided by $3$. All of these examples motivate the study of the asymptotic index of dispersion, $\mathcal{D}$.
@@ 2 Setup and Main Result
Denote the raw rate of births deaths as,
@@
While Theorem 1 is limited to a finite state space, a natural extension is to consider a countably infinite state space where $\lambda_{i}>0$ for all all $i\geq 0$, $\mu_{i}>0$ for all $i\geq 1$, and $q^{\pm}_{i}$ are defined for all $i\geq 0$ with $q_{0}^{-}=0$. In such a case, we also require the standard stability (positive recurrance) condition,
$\sum_{j\in\mathbb{J}}\prod_{k=1}^{j}\frac{\lambda_{k-1}}{\mu_{k}}<\infty.$ (2.12)
In this infinite state space case, $P_{k}$, $\lambda_{k}$, $\Lambda_{k}$, and ${\overline{\lambda}}$ extend naturally as above with,
@@
We say the process is a pure birth-counting process if $q_{i}^{-}=0$ for all states $i$ and $q_{i}^{+}>0$ for at least one state $i$ (note that this term should not be confused with a pure-birth process which is a process with $\mu_{i}=0$ for all $i$). Similarly a pure death-counting process has $q_{i}^{+}=0$ for all state $i$ and $q_{i}^{-}>0$ for at least one state $i$. Further, we call $N(0,t]$ a fully counting pure-birth process if in addition to being pure birth-counting it also has $q_{i}^{+}=1$ for all states $i$. And similarly a fully counting pure-death process is a pure death-counting process that has $q_{i}^{-}=1$ for all states $i$. Finally we call the process fully counting if $q_{i}^{+}=1$ and $q_{i}^{-}=1$ for all states $i$.
Let us see that our finite state space formula, (2.10) from Theorem 1 generalizes the formula, in Theorem 3.1 of [14]. Specifically, [14] studies fully counting pure-death processes, and by switching the role of births and deaths, one may also apply the results to fully counting pure-birth processes. The following corollary applies to this case and one may verify that (2.15) agrees with the results in Theorem 3.1 of [14].
@@
Given a fixed set of birth and death parameters, $\mathcal{D}$ is the same for both the fully counting pure-birth process case, or the fully counting pure-death case. In both of these cases,
@@
Note that in the fully counting pure-birth case, $\Lambda_{k}$ of (2.9) is $\Lambda_{k}^{+}$ of (2.16) and similarly in the fully counting pure-death case, $\Lambda_{k}$ is $\Lambda_{k}^{-}$. Note also that we take $\Lambda_{0}^{-}=0$.
@@
It follows from Equation (2.5) and (2.1) that ${\overline{\lambda}}=\sum_{j=0}^{J-1}\pi_{j}\lambda_{j}$ for both the fully counting pure-birth and fully counting pure-death process. However, considering (2.9) we see that $\Lambda_{k}$ differs for these cases and is $\Lambda_{k}^{-}$ and $\Lambda_{k}^{+}$ for the fully counting pure-death and fully counting pure-birth processes respectively.
Now using Theorem 1, for the fully counting pure-birth case we obtain the following expression for $R_{k}$ in (2.11),
@@
Similarly, for the fully counting pure-death we obtain the same expression on the right hand side. Hence we obtain (2.15) for both of these cases.
@@ 3 Proof of the main result
$\big{(}{\mathbf{W}}^{-1}\big{)}_{ij}=\sum_{k=1}^{\min\{i,j\}}\frac{\pi_{j}}{\pi_{k}\mu_{k}}\qquad(i,j=1,\ldots,J).$ (3.2)
@@
[…]
@@ A Renewal Reward Process
We exploit regeneration epochs in the sample path of the BD process $Q$ at exit times $\tilde{t}$ from the state 0, i.e. $\{Q(\tilde{t}-)=0,Q(\tilde{t})=1\}$, focusing on the generic time $X$ that elapses between such epochs and the number $Y$ of birth and death epochs that are counted (i.e. are not thinned) by the counting function $N(\cdot)$ during this time-interval of length $X$. Then $N(\cdot)$ has asymptotic behaviour the same as a renewal–reward process $C(\cdot)$ for which $\lim_{t\to\infty}t^{-1}\mathop{\rm var}\nolimits C(t)$ is known ([3] and (3.9)–(3.11) below). This derivation shows that the initial distribution of $Q(0)$ plays no part in the theorem when $J$ is finite, i.e. the result depends only on the existence and properties of a stationary distribution $\{\pi_{i}\}$, and thus holds both for a stationary and non-stationary version of the process.
@@ A Transform Based Calculation
To use (3.17), we now find a representation for the moments $n_{1}^{(2)}$, $c_{1}^{(1)}$, and $\tau_{1}^{(2)}$ using first step analysis. Note that this process also yields expressions for the first moments $\tau_{1}^{(1)}$ and $n_{1}^{(1)}$ which naturally agree with those computed above. For such a first step analysis, the r.v.s at (3) are amenable to study via a backwards decomposition in terms of the BD process $Q$ as follows. Given $Q(0)=k$ for some $k=1,\ldots,J$, define $\phi_{k}(s,z)={\mathbb{E}}[{\rm e}^{-s\tau_{k}}z^{N_{J}(\tau_{k})}\mid Q(0)=k]$. The first transition in $Q$ from $Q(0)=k$ occurs after the random time $\sigma_{k}$ which is exponentially distributed with mean $1/r_{k}:=1/(\lambda_{k}+\mu_{k})$, after which $Q=k\pm 1$ (at rate $\lambda_{k}$ for +1, and $\mu_{k}$ for $-1$), and $N_{J}(\tau_{k})$ increases by 1 with probability $q_{k}^{\pm}$ depending on ±$\pm 1$ change in $Q$. Using indicator r.v.s $I(q)=1$ with probability $q$, $=0$ otherwise, these relations can be written as
$(\tau_{k},N_{J}(\tau_{k}))=\Bigg{\{}\begin{aligned} &\big{(}\sigma_{k}+\tau_{k+1},\,I(q_{k}^{+})+N_{J}(\tau_{k+1})\big{)}\;\;\text{with prob. }\lambda_{k}/r_{k},\\ &\big{(}\sigma_{k}+\tau_{k-1},\,I(q_{k}^{-})+N_{J}(\tau_{k-1})\big{)}\;\;\text{with prob. }\mu_{k}/r_{k}.\\ \end{aligned}$ (3.18)
@@
[…]
@@ 4 Examples and Illustrations
We first present examples of finite state queueing systems where there is modeling value for understanding the asymptotic index of dispersion. We then present examples of infinite systems, based on the formal expression in (2.14) with a purpose of exploring the validity of that expression.
@@
BRAVO for the departure process of finite state queues without reneging (M/M/1/K, M/M/s/K, M/M/K/K) was already analyzed extensively in [14] since the main result there (appearing as Corollary 1 in this paper) covers that case. For this current work, we were originally interested in the variance rate of thinned death processes because the output of a queue with reneging has such a structure. That type of process is not covered in [14] since it requires state dependent thinning. We thus have this example:
@@
BRAVO for a queue with reneging. Consider the many-server Poisson system M/M/$s$ with reneging and a finite buffer of size $K$. It is sometimes described as an M/M/$s$/K$+$M system. In terms of a birth–death process on finite state space, we set the state space upper limit $J=K$, the number of servers $s\leq K$, the arrival rate $\lambda>0$, the service rate (per server) $\mu>0$, and the abandonment rate $\gamma\geq 0$. With these, the BD and thinning parameters are,
@@
Setting $\rho=\lambda/(\mu s)$, Figure 1 presents the asymptotic index of dispersion as a function of $\rho$ for various values of $\gamma$. This is for $K=20$, $s=10$ and $\mu=1$. Indeed it is evident that a BRAVO effect appears also in this type of system.
Figure 1: The asymptotic index of dispersion for M/M/$s$/K$+$M systems with $K=20$, $s=10$, and $\mu=1$. When the offered load $\rho\approx\lambda/(\mu s)$, the asymptotic variability of the output process is reduced in comparison to the Poisson process ($\mathcal{D}=1$) case.
This next example, illustrates a situation where counting thinned arrivals maybe of interest:
@@
Finite population interaction upon arrival. Consider a finite population queueing system with $J$ individuals in the population, arriving, and departing from an unlimited service (infinite server) queue. An example is a small population of a certain species living on a small island and occasionally meeting at a fresh water source. As an additional individual arrives to the water source, there is a possibility of it interacting with one of the other individuals already there, or not. This probability increases as the number of individuals at the water source increases.
@@
where we note that $\lambda_{i}$ and $\mu_{i}$ in this form capture the finite population and the unlimited service.
@@
Figure 2: The asymptotic index of dispersion for a counting process of interactions upon arrival to a finite population unlimited service system.
@@
Constant birth rates, fully counting pure-death process. Consider the infinite state space case where $\lambda_{i}=\lambda$ (constant) for all $i$, $q^{+}_{i}=0$ (all $i$), $q_{i}^{-}\equiv 1$, $\mu_{i}=\min\{i,s\}\mu$ and $\rho=\lambda/(s\mu)<1$. This is the case for output of a stable M/M/$s$ queue. By reversibility, [10], it is known that in the stationary case the output process is Poisson and hence has $\mathcal{D}=1$. Indeed using (2.14), ${\overline{\lambda}}=\lambda$ and $\Lambda_{k}=P_{k-1}$ so every product in the numerators in the sum preceding (2.14) is zero, and $\mathcal{D}=1$.

### v2

@@ Abstract
The paper studies the counting process arising as a subset of births and deaths in a birth–death process on a finite state space. Whenever a birth or death occurs, the process is incremented or not depending on the outcome of an independent Bernoulli experiment whose probability is a function of the state of the birth and death process and also depends on whether it is a birth or death that has occurred. We establish a formula for the asymptotic variance rate of this process, also presented as the ratio of the asymptotic variance and the asymptotic mean. Several examples including queueing models illustrate the scope of applicability of the results. An analogous formula for the countably infinite state space is conjectured and tested.
@@ 1 Introduction
We first became aware of the second-order properties discussed in this note in various queueing models in which the simpler cases can be cast purely in a birth–death process setting where computation is more readily effected. In such cases there is a phenomenon, called BRAVO (Balancing Reduces Asymptotic Variance of Outputs), that appears in the limit as the state space grows. For simple output counting processes as in the case of M/M/1/K queues, the study of BRAVO simply requires considering the counting process of deaths in a birth-death process. Yet in more complex models, a more general counting process is required. In this paper we exhibit more general processes $N_{q}$ defined by counting the births and deaths that remain after independent thinnings of these two types of events in a birth–death process and for which the BRAVO characteristics may or may not survive the thinning procedure.
@@
$\mathcal{D}:=\lim_{t\to\infty}\frac{\mathop{\rm var}\nolimits N_{q}(0,t]}{{\mathbb{E}}\big[N_{q}(0,t]\big]},$ (1.3)
namely the limit of the variance of $N_{q}(0,t]$ divided by the expectation of $N_{q}(0,t]$. In mathematical biology, for a general stationary counting process $N_{q}$, $\mathcal{D}$ is known as the Fano factor, as for example in Eden and Kramer [6]. The denominator and numerator here are asymptotically linear in $t\geq 0$, and the limit in (1.3) is always finite and positive. Our main result, Theorem 1, presents a formula for $\mathcal{D}$ in terms of $\{q^{\pm}_{j}:j\in\mathbb{J}\}$ and the birth- and death-rates of $Q$. To our knowledge this condensed explicit formula is new. Yet we mention that since the state space is finite, computable matrix based expressions for $\mathcal{D}$ can be obtained by considering the process as a MAP (Markovian Arrival Process); see for example [9].
In the special case of a pure death-counting process (all $q_{i}^{+}=0$), no thinning ($q^{-}_{i}\equiv 1$), a similar formula was derived in [14] where the BRAVO effect was first noted. BRAVO is a phenomenon observed in certain queueing systems, where if the arrival rate equals (exactly or approximately) the service capacity, then the asymptotic variance of the departure (outputs) process counts reduces in comparison to cases where the arrival rate and service capacity are different. For simple examples of queues, such as the M/M/1/K queue, the study of BRAVO requires study of the process $N_{-}$, as this is the output process of the queue. However, when considering more complex examples, such as queues with reneging (see Example 1 in Section 4), to study the output process we need to introduce state dependent thinning as appearing in the second term of (1.2).
In a sequence of papers associated with BRAVO [4, 5, 7, 8, 11, 12, 13, 14], it was shown that for non-small $J$, it holds that $\mathcal{D}\approx 1$, except for ‘balanced‘ singular cases in which the arrival rate equals the service capacity. In those cases $\mathcal{D}\approx{\cal B}$ where for many standard queueing models, the BRAVO constant ${\cal B}<1$. This constant equals $2/3$ for simple single-server Markovian queues with a finite buffer, [8, 14]; equals $2(1-2/\pi)\approx 0.7268$ for infinite buffer cases with infinite buffers, [1, 13]; and equals $\approx 0.6093$ in many-server queues under the so-called Halfin–Whitt scaling regime, [5]. Note also a correction paper for some of the computations associated with that result, [11].
[…]
@@ 2 Setup and Main Result
Denote the raw rate of births and deaths as,
@@
While Theorem 1 is limited to a finite state space, a natural extension is to consider a countably infinite state space where $\lambda_{i}>0$ for all all $i\geq 0$, $\mu_{i}>0$ for all $i\geq 1$, and $q^{\pm}_{i}$ are defined for all $i\geq 0$ with $q_{0}^{-}=0$. In such a case, we also require the standard stability (positive recurrence) condition,
$\sum_{j\in\mathbb{J}}\prod_{k=1}^{j}\frac{\lambda_{k-1}}{\mu_{k}}<\infty,$ (2.12)
see for example Section 1.3 in [16]. In this infinite state space case, $P_{k}$, $\lambda_{k}$, $\Lambda_{k}$, and $\overline{\lambda}$} extend naturally as above with,
@@
We say the process is a pure birth-counting process if $q_{i}^{-}=0$ for all states $i$ and $q_{i}^{+}>0$ for at least one state $i$ (note that this term should not be confused with a pure-birth process which is a process with $\mu_{i}=0$ for all $i$). Similarly a pure death-counting process has $q_{i}^{+}=0$ for all states $i$ and $q_{i}^{-}>0$ for at least one state $i$. Further, we call $N_{q}(0,t]$ a complete pure birth-counting process if in addition to being pure birth-counting it also has $q_{i}^{+}=1$ for all states $i$. And similarly a complete pure death-counting process is a pure death-counting process that has $q_{i}^{-}=1$ for all states $i$.
Let us see that our finite state space formula, (2.10) from Theorem 1 generalizes the formula, in Theorem 3.1 of [14]. Specifically, [14] studies complete pure death-counting processes, and by switching the role of births and deaths, one may also apply the results to complete pure birth-counting processes. The following corollary applies to this case and one may verify that (2.15) agrees with the results in Theorem 3.1 of [14].
@@
Given a fixed set of birth and death parameters, $\mathcal{D}$ is the same for both the complete pure birth-counting process case, or the complete pure death-counting process case. In both of these cases,
@@
Note that in the complete pure birth-counting process case, $\Lambda_{k}$ of (2.9) is $\Lambda_{k}^{+}$ of (2.16) and similarly in the complete pure death-counting process case, $\Lambda_{k}$ is $\Lambda_{k}^{-}$. Note also that we take $\Lambda_{0}^{-}=0$.
@@
It follows from Equation (2.5) and (2.1) that ${\overline{\lambda}}=\sum_{j=0}^{J-1}\pi_{j}\lambda_{j}$ for both the complete pure birth-counting process and complete pure death-counting process. However, considering (2.9) we see that $\Lambda_{k}$ differs for these cases and is $\Lambda_{k}^{-}$ and $\Lambda_{k}^{+}$ for the complete pure birth-counting process and complete pure death-counting process respectively.
Now using Theorem 1, for the complete pure birth-counting process case we obtain the following expression for $R_{k}$ in (2.11),
@@
Similarly, for the complete pure death-counting process we obtain the same expression on the right hand side. Hence we obtain (2.15) for both of these cases.
@@ 3 Proof of the main result
$\big({\mathbf{W}}^{-1}\big)_{ij}=\sum_{k=1}^{\min\{i,j\}}\frac{\pi_{j}}{\pi_{k}\mu_{k}}\qquad(i,j=1,\ldots,J),$ (3.2)
where $\pi_{i}$ is given by (2.2).
@@
$\qquad\qquad\qquad+\frac{\pi_{i}}{\pi_{i}\mu_{i}}\Big[(\lambda_{i}+\mu_{i})-\frac{\pi_{i+1}}{\pi_{i}}\mu_{i+1}\Big]\,=\,0+1\,=\,1.$
@@ A Renewal Reward Process
We exploit regeneration epochs in the sample path of the BD process $Q$ at exit times $\tilde{t}$ from the state 0, i.e. $\{Q(\tilde{t}-)=0,Q(\tilde{t})=1\}$, focusing on the generic time $X$ that elapses between such epochs and the number $Y$ of birth and death epochs that are counted (i.e. are not thinned) by the counting function $N_{q}(\cdot)$ during this time-interval of length $X$. Then $N_{q}(\cdot)$ has asymptotic behaviour the same as a renewal–reward process $C(\cdot)$ for which $\lim_{t\to\infty}t^{-1}\mathop{\rm var}\nolimits C(t)$ is known ([3] and (3.9)–(3.11) below). This derivation shows that the initial distribution of $Q(0)$ plays no part in the theorem when $J$ is finite, i.e. the result depends only on the existence and properties of a stationary distribution $\{\pi_{i}\}$, and thus holds both for a stationary and non-stationary version of the process.
@@ A Transform Based Calculation
To use (3.17), we now find a representation for the moments $n_{1}^{(2)}$, $c_{1}^{(1)}$, and $\tau_{1}^{(2)}$ using first step analysis. Note that this process also yields expressions for the first moments $\tau_{1}^{(1)}$ and $n_{1}^{(1)}$ which naturally agree with those computed above. For such a first step analysis, the r.v.s at (3) are amenable to study via a backwards decomposition in terms of the BD process $Q$ as follows. Given $Q(0)=k$ for some $k=1,\ldots,J$, define $\phi_{k}(s,z)={\mathbb{E}}[{\rm e}^{-s\tau_{k}}z^{N_{J}(\tau_{k})}\mid Q(0)=k]$. The first transition in $Q$ from $Q(0)=k$ occurs after the random time $\sigma_{k}$ which is exponentially distributed with mean $1/r_{k}:=1/(\lambda_{k}+\mu_{k})$, after which $Q=k\pm 1$ (at rate $\lambda_{k}$ for +1, and $\mu_{k}$ for 1-1), and $N_{J}(\tau_{k})$ increases by 1 with probability $q_{k}^{\pm}$ depending on $\pm 1$ change in $Q$. Using indicator r.v.s $I(q)=1$ with probability $q$, =0=0 otherwise, these relations can be written as
[…]
@@
The notation ${\mathbf{a}}\circ{\mathbf{b}}$ is useful and has $({\mathbf{a}}\circ{\mathbf{b}})\circ{\mathbf{c}}={\mathbf{a}}\circ({\mathbf{b}}\circ{\mathbf{c}})={\mathbf{a}}\circ{\mathbf{b}}\circ{\mathbf{c}}$ and $\boldsymbol{\pi}^{\top}({\mathbf{a}}\circ{\mathbf{b}})={\bf 1}^{\top}(\boldsymbol{\pi}\circ{\mathbf{a}}\circ{\mathbf{b}})$, but has the drawback that, for a general $J\times J$ matrix $\mathbf{A}$}, ${\mathbf{A}}({\mathbf{a}}\circ{\mathbf{b}})\neq({\mathbf{A}}{\mathbf{a}})\circ{\mathbf{b}}$. We also define the $J\times J$-matrix ${\mathbf{V}}=:(v_{ij})$ for $i,j\in\{1,\ldots,J\}$ with $v_{ij}=0$ except for its super- and sub-diagonal elements $v_{i,i+1}=\lambda_{i}q_{i}^{+}$ ($i=1,\ldots,J-1$) and $v_{i,i-1}=\mu_{i}q^{-}_{i}$ $(i=2,\ldots,J)$. This matrix helps facilitate algebraic manipulation.
@@ 4 Examples and Illustrations
We first present examples of finite state queueing systems where there is modeling value for understanding the asymptotic index of dispersion. We then present examples of infinite systems, based on the formal expression in (2.14) with a purpose of exploring the validity of that expression. Note that not all examples are related to BRAVO. In particular, in some of the examples below, the process $N_{q}$ is not related to the output counting process of a queueing system.
@@
Let us consider two examples arising from finite state queueing models. The first example illustrates the existence of a BRAVO phenomenon in queues with reneging, while the second example studies a process not related to the output of a queue or the BRAVO effect. Note that the departure process of finite state queues without reneging (M/M/1/K, M/M/s/K, M/M/K/K) was already analyzed extensively in [14] since the main result there (appearing as Corollary 1 in this paper) covers that case. For such queues, BRAVO was observed. For this current work, we were originally interested in the variance rate of thinned death processes because the output of a queue with reneging has such a structure. That type of process is not covered in [14] since it requires state dependent thinning.
Figure 1: The asymptotic index of dispersion for M/M/ss/K++M systems with $K=20$, $s=10$, and $\mu=1$. When the offered load $\rho\approx 1$, the asymptotic variability of the output process is reduced in comparison to the Poisson process ($\mathcal{D}=1$) case.
@@
BRAVO for a queue with reneging. Consider the many-server Poisson system M/M/$s$ with reneging and a finite buffer of size $K$. It is sometimes described as an M/M/ss/K++M system. In terms of a birth–death process on finite state space, we set the state space upper limit $J=K$, the number of servers $s\leq K$, the arrival rate $\lambda>0$, the service rate (per server) $\mu>0$, and the abandonment rate $\gamma\geq 0$. With these, the BD and thinning parameters are,
@@
Setting $\rho=\lambda/(\mu s)$, Figure 1 presents the asymptotic index of dispersion as a function of $\rho$ for various values of $\gamma$. This is for $K=20$, $s=10$ and $\mu=1$. Indeed it is evident that a BRAVO effect appears also in this type of system. It is evident that for $\gamma=0$ the minimizer is at $\rho=1$ (this case was also covered in [14]), whereas for $\gamma>0$ the minimizer deviates slightly from $\rho=1$. There is room for an asymptotic analysis here, similar in nature to [5], considering the effects of $\gamma$, as well as $K\to\infty$ and $s\to\infty$. We leave such an asymptotic analysis for further research, yet observe at this point that indeed a form of the BRAVO effect holds.
@@
Finite population interaction upon arrival. This example illustrates a situation where counting thinned arrivals (as opposed to departures) maybe of interest. We note that this example is not about BRAVO. In particular, this example does not illustrate existence/non-existence of a BRAVO effect since it does not focus on the departure process of a queue. Nevertheless, the resulting curve of $\mathcal{D}$ exhibits interesting behavior. Note that we informally refer to this example as “counting animal fights in a billabong”.
Consider a finite population queueing system with $J$ individuals in the population, arriving to, and departing from an unlimited service (infinite server) queue (billabong). An example is a small finite population of a certain species occasionally meeting at a fresh water source (billabong). As an additional individual arrives to the water source, there is a possibility of it interacting (e.g. fighting) with one of the other individuals already there, or not. This probability increases as the number of individuals at the water source increases. With this, we are interested in the counting process of the number of interactions (fights) occurring at the water source.
@@
where we note that $\lambda_{i}$ and $\mu_{i}$ in this form capture the finite population and the unlimited service. Further, $q_{i}^{+}$ captures the increasing probability of an interaction (fight) upon arrival.
@@
Figure 2: The asymptotic index of dispersion for a counting process of interactions upon arrival to a finite population infinite-service system.
@@
Constant birth rates, complete pure death-counting process. Consider the infinite state space case where $\lambda_{i}=\lambda$ (constant) for all $i$, $q^{+}_{i}=0$ (all $i$), $q_{i}^{-}\equiv 1$, $\mu_{i}=\min\{i,s\}\mu$ and $\rho=\lambda/(s\mu)<1$. This is the case for output of a stable M/M/$s$ queue. By reversibility, [10], it is known that in the stationary case the output process is Poisson and hence has $\mathcal{D}=1$. Indeed using (2.14), ${\overline{\lambda}}=\lambda$ and $\Lambda_{k}=P_{k-1}$ so every product in the numerators in the sum preceding (2.14) is zero, and $\mathcal{D}=1$.